OpenStudy (yolo_king):

The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (-3, 2)

OpenStudy (yolo_king):

A) y = 5x + 13 B)\[y=\frac{ -1 }{ 5 }x+\frac{ 7 }{ 5 }\] C)\[y=-\frac{ 1 }{ 5 }x -\frac{ 7 }{ 5 }\] D)y = 5x - 17

OpenStudy (yolo_king):

@iGreen.

OpenStudy (yolo_king):

@iGreen. MY SAVIOuR HELP ME

OpenStudy (anonymous):

Perpendicular means it has opposite slope. The slope here is 5, so what's the opposite of that?

OpenStudy (yolo_king):

-5 qq

OpenStudy (anonymous):

\(5 \rightarrow\dfrac{5}{1}\) 5 is the same thing as 5/1, so flip 5/1 and make it negative.

OpenStudy (anonymous):

No, -5 is incorrect.

OpenStudy (yolo_king):

Oh right, right

OpenStudy (yolo_king):

So would the answer be c

OpenStudy (anonymous):

Hold on a minute.

OpenStudy (anonymous):

The slope is correct, it should be \(-\dfrac{1}{5}\)

OpenStudy (anonymous):

Actually, no C is incorrect.

OpenStudy (yolo_king):

Hmmmmm.... let me redo dis

OpenStudy (yolo_king):

A is the slope of the perpendicular line its \[-\frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

Yes, we should plug it in point-slope form. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. So we have: \(y - 2 = -\dfrac{1}{5}(x - (-3))\) or \(y - 2 = -\dfrac{1}{5}(x + 3)\) Distribute -1/5 into the parnethesis: \(y - 2 = -\dfrac{1}{5}x - \dfrac{3}{5}\) Now add 2 to both sides, what's -3/5 + 2?

OpenStudy (yolo_king):

Ermmmm.... 7/5<----answer... i think

OpenStudy (anonymous):

Nvm, yes, 7/5 so we have: \(y = -\dfrac{1}{5}x + \dfrac{7}{5}\)

OpenStudy (yolo_king):

Oh I thought the 3 was only negative heh my mistake

OpenStudy (anonymous):

Lol, that's my mistake :P

OpenStudy (anonymous):

Good work

OpenStudy (yolo_king):

YAy C:

OpenStudy (yolo_king):

Gg

OpenStudy (yolo_king):

Thanks man!

OpenStudy (yolo_king):

So would the answer be B right just for clarification

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (yolo_king):

Alright